OpenStudy (anonymous):

When rolling a number cube, what is the probability of rolling a number greater than 2?

OpenStudy (anonymous):

okay so there are 6 numbers you can roll, and you want to know what the probability of rolling higher than 2 is. well the odds are 1/6 for rolling any number and there are 4 numbers greater than 2 on a number cube so your probability is 1/6 + 1/6 + 1/6 + 1/6 = 4/6 = 2/3

OpenStudy (anonymous):

thank you lots :)

OpenStudy (anonymous):

1/3 is not one of the given choices!

OpenStudy (anonymous):

the answer was 2/3(: